graph of p vs 1/v square boyles law
plz tell the ans and how to do it?

Grade:11

1 Answers

Japneet Kaur
3 years ago
to plot P vs 1/V2 graph
P=y
1/V=x or V=1/\sqrt{x}
Accoding to Boyles Law, 
PV= constant
y/\sqrt{x} = K
y= Kx2 , which is the equation of parabola